OpenStudy (anonymous):

Can someone please help me? Given the triangle and equation below, what is the value of x? csc∠A + tan∠C = 5x

OpenStudy (anonymous):

Here is the attachment And the choices

OpenStudy (raden):

first, find AB with pythagorean therem

OpenStudy (anonymous):

ok

OpenStudy (raden):

got it ?

OpenStudy (raden):

hint : AB^2 = AC^2 - BC^2 AB^2 = 7^2 - 6^2 solve for AB

OpenStudy (anonymous):

Is the 2nd answer correct?

OpenStudy (tkhunny):

One day, you'll have to start showing your work and actually demonstrating some competence in the material. Please stop guessing.